# DetailCLIP: Detail-Oriented CLIP for Fine-Grained Tasks

### v1

@@ 3.2.2 Pixel-Level Reconstruction
DeiT[42] IN-1K 300 720M 47.4 - 44.1 39.8
MAE[16] IN-1K 800 960M 46.5 34.3 46.2 39.1
DINO[5] IN-1K 800 9600M 46.8 34.5 47.4 40.1
iBOT[56] IN-1K 300 720M 47.3 34.7 48.4 42.1
AttMask[22] IN-1K 300 4320M 47.5 35.2 48.9 42.2
@@
CLIP[38] 400M - - 46.4 34.2 43.6 39.5
SLIP[33] YFCC-15M 25 750M 46.6 36.1 44.0 40.3
MaskCLIP[12] YFCC-15M 25 750M 47.5 36.3 45.4 40.9
A-CLIP[48] YFCC-15M 25 750M 47.0 34.7 45.8 41.7
@@
Table 1: Performance comparison of various models on detail-oriented visual tasks, including segmentation and object detection. All models utilize the vision component’s ViT-B (Vision Transformer Base) architecture. Segmentation results are presented for the ADE20K dataset using UperNet and linear decoders. At the same time, object detection performance is evaluated on the MS COCO dataset with Average Precision for the bounding box. DetailCLIP consistently outperforms other models, particularly in tasks requiring high levels of detail and precision. (”Effective View” is the total number of image views used during training, calculated as the number of views per image multiplied by the number of epochs and size of the dataset. This metric is critical to our evaluation as it reflects the model’s exposure to diverse perspectives of the training data, which directly impacts its ability to generalize and perform well on detail-oriented tasks.)
3.2.3 Token Removal
Token removal in image models, particularly in ViT, is a strategy to refine the attention mechanism by selectively eliminating tokens that contribute less to the final decision-making process. The token removal method allows the model to dynamically down-weight or ignore less important patches, thereby reducing computational load and focusing attention on the most relevant areas of the image, ultimately enhancing both performance and efficiency.
Various approaches for token removal have been proposed, such as random token removal [22] and attentive token removal [48]. In our work, we introduce a novel technique that considers the specific functions of our model, including textual information processing, patch comparison, and image reconstruction. As demonstrated in 4.2, our token removal strategy outperforms existing methods by addressing both fine-grained and coarse-grained details simultaneously.
As illustrated in Figure 2, our token removal process operates as follows. We first input an image view into the teacher encoder model ($g_{\theta}$). The attention values generated by the teacher encoder are then used to mask the $50\%$ of patches with the lowest values. Specifically, for each patch $p_{\text{t}}$ the attention values are calc ulated using the following formula:
  • $N$ is the total number of attention heads across all layers.
  • $Q_{i}$ is the query of the [CLS] from the $i$-th attention head.
  • $K_{i}(p_{t})$ is the key of the patch at position $p_{\text{t}}$ from the $i$-th attention head.
  • $d$ is the dimensionality of the query and key.
As shown in Figure 1, masking image patches by removing the lowest half of the attention values (AV) calculated by the teacher encoder results in the images displayed in the figure.
3.2.4 Integrated Loss Function
In this section, we introduce the formulation of a comprehensive loss function that combines the various loss components essential for effective multi-task learning. The integrated loss function is defined as:
In this equation, the hyperparameters $\alpha_{1}$, $\alpha_{2}$, and $\alpha_{3}$ determine the relative importance of each loss term. These weights are set to $1$ in our experiments unless specified otherwise. The comprehensive loss function is designed to enable the model to concurrently learn across the four tasks introduced in previous sections, ensuring a balanced and integrated training process.
Additionally, in the ablation study, we explore the impact of different hyperparameter choices, demonstrating how varying these weights can influence the overall model performance and task-specific outcomes.
4 Experiments
4.1 Setup
Our computational setup is designed to support the extensive experiments conducted in this study. It consists of four nodes, each equipped with four NVIDIA A100 GPUs, each with 80GB.
4.1.1 Training Data and Augmentation Strategy
Our model is trained on a carefully curated 15 million image subset [38] of the YFCC100M dataset [40], which contains only English-language titles and descriptions. For each image, a valid caption—either a title or description—is randomly selected during training, following the methodology used in SLIP [33].
To enhance model robustness, we employ a data augmentation strategy similar to that of SLIP. Images are randomly resized and cropped to a size between $50\%$ and $100\%$ of their original dimensions. This augmentation is applied to the images in the online training branches, allowing the model to learn from varied perspectives of the same image.
For the teacher part, a slightly different approach is taken. We use a larger randomly cropped sub-image compared to the online views, which allows us to accurately compute attention value.
@@ 4.1.1 Training Data and Augmentation Strategy
Table 2: Performance comparison of various models on text-image retrieval tasks across the Flickr30K and MS COCO datasets, as well as zero-shot evaluation on ImageNet-1K (IN-1K). The table reports results after 25 and 50 epochs ($E$), showing performance metrics for image-to-text (I2T) and text-to-image (T2I) retrieval on each dataset. The results were collected in the A-CLIP paper [48].
4.1.2 Architecture and Training Setting
To ensure a fair comparison with existing models, our visual encoders ($g_{\theta}$ and $g_{\phi}$) are based on the widely recognized Vision Transformer (ViT-B/16) [13] architecture. This model features $12$ layers, each with a width of $768$ and $12$ attention heads.
For the text encoder ($e_{\gamma}$), we adopted a 12-layer Transformer architecture with a width of $512$ and $8$ attention heads, following the design principles established by the CLIP model [38].
The decoder ($d_{\phi}$) itself is composed of a series of transformer blocks like MAE paper [16]. The decoder handles the entire set of input patches, including both the attention-masked and retained patches. Positional embeddings are added to each token in the set to preserve the spatial information.
The projection heads ($h_{\theta}$ and $h_{\phi}$) in our framework are implemented as 3-layer MLPs, featuring an L2-normalized bottleneck, similar to the approach used in DINO [5]. The output dimension of the shared projection head is set to $8192$ like iBOT paper [56], for robust feature representation and alignment during training.
For our experiments, we employ the $AdamW$ optimizer [30] with a learning rate of $45e-$ 4 and a weight decay of $0.5$. The training is conducted with a substantial batch size of $4096$.
4.2 Experimental Analysis
@@ Methods
Table 3: Zero-shot performance evaluation across multiple classification benchmarks. Epochs indicate the total number of training epochs. The results were collected in the A-CLIP paper [48].
4.2.1 Detail-Oriented Visual Tasks
In order to comprehensively evaluate the effectiveness of our proposed DetailCLIP framework, we designed a series of experiments focused on detail-oriented visual tasks. Specifically, we implemented different distinct tasks that emphasize the model’s ability to capture fine-grained details in complex visual environments.
Semantic Segmentation on ADE20K: To evaluate the segmentation capabilities of DetailCLIP, we conducted a series of experiments on the ADE20K dataset [54] using different decoding strategies. First, we tested the model with the UperNet [45] decoder, a well-established architecture for semantic segmentation, to assess its ability to delineate object boundaries and achieve high segmentation accuracy. Next, we employed a linear decoder for segmentation tasks to explore the model’s versatility. This more straightforward approach allowed us to evaluate how well DetailCLIP maintains performance with minimal architectural complexity. These experiments collectively provide a comprehensive understanding of DetailCLIP’s strengths in detail-oriented segmentation tasks.
We utilized the UperNet or linear decoder with an input resolution of $512\times 512$ pixels, conducting end-to-end training for $160k$ iterations. The performance of our model was evaluated using the mean Intersection over Union (mIoU) metric.
Object Detection and Instance Segmentation on the COCO: This task requires simultaneous object localization and classification, making them complex tasks that demand high precision. For our evaluation, we employed the Cascade Mask R-CNN model[4, 18] like iBOT [56], known for effectively producing bounding boxes and instance masks on the COCO dataset [27]. The evaluation metrics used in this study were the Average Precision for bounding boxes ($AP^{b}$) and masks ($AP^{m}$).
Evaluating Task Outcomes: Table 1 highlights the significant advantages of DetailCLIP over existing models in various detail-oriented visual tasks. Utilizing the ViT-B architecture, DetailCLIP outperformed other models across multiple benchmarks. For instance, on the ADE20K dataset with the UperNet decoder, DetailCLIP achieved a $mIoU$ of $48.8$, $1.3$ points higher than the closest competitor, MaskCLIP, which recorded a $mIoU$ of $47.5$. Similarly, in the linear decoder setting, DetailCLIP’s $mIoU$ of $39.3$ surpasses SLIP’s performance by $3.2$ points, highlighting our model’s superior capacity to maintain high performance.
Furthermore, in object detection tasks on the MS COCO dataset, DetailCLIP for bounding boxes achieved $48.9$ $AP^{b}$, outperforming previous models by as much as $3.1$ points, with ACLIP being the closest at $45.8$. The mask precision is $42.5$ $AP^{m}$, further underscoring our model’s effectiveness, exceeding the best-performing alternative by $0.8$ points. These substantial improvements demonstrate DetailCLIP’s effectiveness in enhancing high-level semantic understanding and fine-grained detail extraction, setting a new benchmark in the field.
Visual Comparative Analysis: Figure 5 shows the result of our model compared to the other baselines. We invite you to refer to the appendix. This comparison covers segmentation and object detection tasks, demonstrating DetailCLIP’s enhanced performance.
Image Ground Truth A-Clip SLIP DetailCLIP
4.2.2 Image Classification
Zero-shot on Text-Image Retrieval: We also present the zero-shot text-image retrieval on 3 benchmark datasets: Flickr30K [50], MS-COCO [27], and ImageNet-1K [11]. Our findings indicate that using plain text, without any added prefixes or suffixes, consistently yields strong performance across all models evaluated.
Table 2 comprehensively evaluates different models on text-image retrieval tasks, explicitly focusing on Flickr30K and MS-COCO datasets, also a zero-shot evaluation on ImageNet-1K.
Zero-Shot Classification Performance Evaluation on Diverse Benchmarks: We evaluated the zero-shot classification performance of our proposed DetailCLIP approach across $13$ different classification tasks, following the evaluation protocols established by SLIP [33]. This included using identical prompts during testing to ensure a fair comparison. This rigorous evaluation framework allowed us to directly assess and compare DetailCLIP’s performance against other leading models.
The results in the table 3 demonstrate that DetailCLIP consistently outperforms its competitors. Specifically, DetailCLIP achieved the highest performance in $12$ out of $26$ evaluation scenarios and secured the second-best position in $5$ additional cases. Notably, DetailCLIP also achieved the highest average scores in both the 25-epoch ($37.3$) and 50-epoch ($37.6$) settings, underscoring its robustness and generalization capabilities across a diverse array of benchmarks.
These results highlight DetailCLIP’s effectiveness in zero-shot classification tasks, particularly compared to other state-of-the-art models across various datasets.
@@ 4.2.3 Ablation Study on Loss Weight Optimization
5 Conclusion
In this paper, we introduced DetailCLIP, a novel framework designed to overcome traditional CLIP-based models’ limitations in fine-grained and coarse-grained tasks. While existing models like CLIP excel at globally aligning image and text representations, they often fail to capture the fine-grained details necessary for tasks such as image segmentation and object detection. To address this, we integrated innovative techniques into DetailCLIP, including patch-level comparison, pixel-level reconstruction, and an attention-based token removal mechanism. These features enable the model to focus on the most critical regions of an image, significantly enhancing its performance in detail-oriented tasks. Our extensive experiments demonstrated that DetailCLIP consistently outperforms state-of-the-art models in tasks requiring high levels of detail, such as segmentation on the ADE20K dataset and object detection on the MS COCO dataset, where it sets new benchmarks in critical metrics like mIoU and Average Precision. Furthermore, DetailCLIP’s strength extends beyond fine-grained tasks, as it excels in coarse-grained tasks like zero-shot classification.
Appendix A Visual Analysis

### v2

@@ 4.2.1 Detail-Oriented Visual Tasks
DeiT IN-1K 300 720M 47.4 - 44.1 39.8
MAE IN-1K 800 960M 46.5 34.3 46.2 39.1
DINO IN-1K 800 9600M 46.8 34.5 47.4 40.1
iBOT IN-1K 300 720M 47.3 34.7 48.4 42.1
AttMask IN-1K 300 4320M 47.5 35.2 48.9 42.2
@@
CLIP 400M - - 46.4 34.2 43.6 39.5
SLIP YFCC-15M 25 750M 46.6 36.1 44.0 40.3
MaskCLIP YFCC-15M 25 750M 47.5 36.3 45.4 40.9
A-CLIP YFCC-15M 25 750M 47.0 34.7 45.8 41.7
@@
Table 1: Performance comparison on segmentation (ADE20K) and object detection (MS COCO) tasks using the ViT-B architecture. Segmentation results are based on UperNet and linear decoders, while object detection uses bounding box Average Precision. “Effective View” indicates the model’s training exposure, calculated as views per image × epochs × dataset size.
We evaluated DetailCLIP on detail-oriented visual tasks to test its ability to capture fine-grained details in complex visual environments.
Semantic Segmentation on ADE20K: We evaluated DetailCLIP on the ADE20K dataset (Zhou et al., 2017) using both UperNet (Xiao et al., 2018) and linear decoders. UperNet assessed the model’s ability to capture detailed object boundaries, while the linear decoder tested performance under minimal architectural complexity. Input images had a resolution of $512\times 512$ pixels, and training was conducted for $160k$ iterations. Performance was measured using the mean Intersection over Union (mIoU) metric.
Object Detection and Instance Segmentation on COCO: We evaluated DetailCLIP using the Cascade Mask R-CNN model (Cai & Vasconcelos, 2019; He et al., 2017), following iBOT, to generate bounding boxes and instance masks on the COCO dataset (Lin et al., 2014). Performance was measured using Average Precision for bounding boxes ($AP^{b}$) and masks ($AP^{m}$).
Evaluating Task Outcomes: Table 1 shows DetailCLIP’s strong performance across detail-oriented visual tasks. Using the ViT-B architecture, DetailCLIP achieved a $mIoU$ of $48.8$ on ADE20K with UperNet, surpassing MaskCLIP by $1.3$ points. In the linear decoder setting, DetailCLIP’s $mIoU$ of $39.3$ outperformed SLIP by $3.2$ points. In object detection on MS COCO, DetailCLIP achieved $48.9$ $AP^{b}$ for bounding boxes, surpassing ACLIP by $3.1$ points, and $42.5$ $AP^{m}$ for masks, $0.8$ points higher than the best alternative. These improvements highlight DetailCLIP’s effectiveness in capturing both high-level semantics and fine-grained details.
Visual Comparative Analysis: Figure 2 shows the result of our model compared to the other baselines. Additionally, we invite you to refer to Section A.5 of the appendix, where we cover segmentation and object detection tasks.
Image Ground Truth A-Clip SLIP DetailCLIP
4.2.2 Image Classification
Zero-shot on Text-Image Retrieval: We also present the zero-shot text-image retrieval on 3 benchmark datasets: Flickr30K (Young et al., 2014), MS-COCO (Lin et al., 2014), and ImageNet-1K (Deng et al., 2009). Our findings indicate that using plain text, without any added prefixes or suffixes, consistently yields strong performance across all models evaluated. Table 2 comprehensively evaluates different models on text-image retrieval tasks, explicitly focusing on Flickr30K and MS-COCO datasets, also a zero-shot evaluation on ImageNet-1K.
Additionally, in Section A.4 of the appendix, we also provide the effect of different hyperparameter settings on ImageNet-1K’s zero-shot scores.
Zero-Shot Classification Performance Evaluation on Diverse Benchmarks: We evaluated DetailCLIP’s zero-shot classification performance on 13 tasks, following SLIP’s (Mu et al., 2022) evaluation protocols with identical prompts for fair comparison. The results in the table 3 demonstrate that DetailCLIP consistently outperforms its competitors. Specifically, DetailCLIP achieved the highest performance in $12$ out of $26$ evaluation scenarios and secured the second-best position in $5$ additional cases. DetailCLIP also achieved the highest average scores in both the 25-epoch ($37.3$) and 50-epoch ($37.6$) settings.
@@ 4.2.2 Image Classification
Table 2: Comparison of model performance on text-image retrieval for Flickr30K, MS COCO, and zero-shot ImageNet-1K (IN-1K). Metrics for image-to-text (I2T) and text-to-image (T2I) retrieval are shown after 25 and 50 epochs ($E$), based on results from A-CLIP Yang et al. (2023).
@@ Methods
Table 3: Zero-shot performance evaluation across multiple classification benchmarks. Epochs indicate the total number of training epochs. The results were collected in the A-CLIP paper Yang et al. (2023).
5 Conclusion
In this paper, we introduce DetailCLIP, a self-improving vision-language framework designed to overcome the limitations of traditional CLIP-based models in both fine- and coarse-grained tasks. While CLIP excels at aligning global image-text representations, it often misses the fine-grained details needed for tasks like segmentation and object detection. To address this, DetailCLIP incorporates self-curated learning objectives—patch-level self-distillation, pixel-level reconstruction, and attention-based token filtering—to iteratively refine both high-level semantics and detailed features. Experiments show that DetailCLIP generalizes well across tasks, surpassing state-of-the-art models on detail-focused benchmarks like ADE20K and MS COCO, while also delivering strong performance on coarse-grained tasks such as zero-shot classification.
Acknowledgments
This project was made possible, in part, by support from the National Science Foundation grant number OAC-2018627 and from the Ohio Supercomputer Center.
Appendix A Appendix
A.1 Attention Token Removal
Figure 3: Illustration of the token removal mechanism, which retains semantically significant tokens to focus on critical input aspects while discarding unnecessary details.
In Figure 3, the mechanism of token removal is illustrated. The process selectively eliminates less relevant tokens, retaining only those with strong semantic significance. This approach enhances the model’s focus on the most important aspects of the input, ensuring that critical information is preserved while unnecessary details are discarded. The current settings of the model use a 50% mask ratio, preserving only the semantically relevant details of the image. For example, in the image with the two dogs, the grass background is removed as it is less important to the overall image. Similarly, in the image of the black bird, the tree and branches are ignored while preserving the bird, as it is central to the image. In the DetailCLIP model, the attention scores are derived from the teacher network and are used to guide the student network during training.
A.2 Computational Efficiency
Our model demonstrates an effective balance between performance and efficiency, as highlighted in Table 4 and Table 2. It requires only 24G of GPU memory, significantly less than SLIP (30G), while achieving a training time of 1.16x—closely aligned with CLIP (1.00x) and notably faster than SLIP (2.67x) and MaskCLIP (1.56x). Our resource usage and training speed make DetailCLIP highly suitable for applications with limited GPU resources, while offering strong performance on downstream tasks.
Methods Training Time GPU Memory
CLIP 1.00x 14G
SLIP 2.67x 30G
MaskCLIP 1.56x 16G
A-CLIP 1.16x 14G
DetailCLIP 1.16x 24G
Table 4: Training time and GPU memory usage of various models, measured consistently on the same device. Training time is reported relative to the baseline CLIP (1.00x), and GPU memory is shown in gigabytes (GB) Yang et al. (2023).
A.3 Pseudocode
Algorithm 1 presents the pseudocode for the DetailCLIP algorithm, which integrates CLIP, self-distillation, and pixel-level loss. The steps include updating the momentum encoder using EMA, generating teacher embeddings for both views (without gradients), and creating attention-based masks with a 50% mask ratio. Student embeddings are then obtained after applying the masks and projected using CLIP projections alongside text embeddings. The model produces outputs for student distillation and reconstruction, followed by computing classification, patch-level, and reconstruction losses to calculate the total loss.
Algorithm 1 Pytorch-like code for DetailCLIP
# u, v: view 1 and view 2 of the image
# s_visual, t_visual: student and teacher image encoders
# t_dist_head, s_dist_head: teacher and student distillation heads
# encode_image, encode_text: students CLIP image and text projections
# reconstr_head: pixel-level reconstruction head
def forward(u, v, text, momentum):
# Update momentum encoder using EMA
update_momentum_encoder(momentum)
# Teacher embeddings for both views (no gradient)
with no_grad():
t_embed_u, attn_u = t_visual(u)
t_embed_v, attn_v = t_visual(v)
t_dist_out = t_dist_head(concat(t_embed_u, t_embed_v))
# Generate masks from attention using 50% mask ratio
mask_u, mask_v = get_att_mask(attn_u), get_att_mask(attn_v)
# Get student embeddings after applying masks
s_embed_u = s_visual(u, mask=mask_u)
s_embed_v = s_visual(v, mask=mask_v)
# Project student embeddings using CLIP projections
img_emb_u = encode_image(s_embed_u)
img_emb_v = encode_image(s_embed_v)
# Get text embedding
text_emb = encode_text(tokenize(text))
# Student distillation output
s_dist_out = s_dist_head(concat(s_embed_u, s_embed_v))
# Reconstruction
s_dec_u = decoder(s_embed_u)
s_dec_v = decoder(s_embed_v)
u_rec = h_reconst(s_dec_u)
v_rec = h_reconst(s_dec_v)
# Compute losses
clip_loss = compute_clip_loss(img_emb_u, img_emb_v, text_emb)
distill_loss = compute_distill_loss(s_dist_out, t_dist_out)
recon_loss = compute_recon_loss(u_rec, v_rec, mask_u, mask_v)
# Total loss
total_loss = clip_loss + distill_loss + recon_loss
return total_loss
A.4 Ablation Study on Loss Weight
@@
A.5 Visual Analysis